# Learning not cheating: AI assistance can enhance rather than hinder skill development

### v2

@@ Abstract
It is widely believed that outsourcing cognitive work to AI boosts immediate productivity at the expense of long-term human capital development. An overlooked possibility is that AI tools can support skill development by providing just-in-time, high-quality, personalized examples. In this investigation, lay forecasters predicted that practicing writing cover letters with an AI tool would impair learning compared to practicing writing letters without the tool. However, in a highly-powered pre-registered experiment, participants randomly assigned to practice writing with AI improved more on a writing test one day later compared to writers assigned to practice without AI. Notably, writers given access to the AI tool improved more despite exerting less effort, whether measured by time on task, keystrokes, or subjective ratings. We replicated and extended these results in a second pre-registered experiment, showing that writers given access to the AI tool again outperformed those who practiced on their own—but performed no better than writers merely shown an AI-generated cover letter that they could not edit. Collectively, these findings constitute an existence proof that by providing personalized examples of high-quality work, AI tools can improve, rather than undermine, learning.
@@
Nevertheless, there is growing concern that AI tools will be used as a crutch, providing immediate gains in performance at the expense of long-run development of human capital (5, 6, 7). For instance, in a 2024 poll, 62% of surveyed adults predicted that Generative AI will “lead to humans becoming less intelligent” (8). In January 2023, New York City public schools banned ChatGPT, citing “concerns about negative impacts on student learning (9).” When this ban was lifted three months later, it was not because of the potential of AI to scaffold learning, but instead because of the “reality that students are participating in and will work in a world where understanding Generative AI is crucial (10).” The sentiment behind the initial ban aligns with teacher perceptions: in a nationally representative poll in May 2024, four times as many K-12 educators judged the use of AI tools as net harmful (24%) than net beneficial (6%) (11).
@@
Second, regardless of accuracy, the fluent and instantaneous solutions AI tools generate may contribute to an illusion of mastery. To the extent users conflate the skills of an AI tool with their own, they may be less likely to seek feedback and improve. Prior research has found that searching for information on the Internet, for example, creates an illusion whereby people conflate knowledge outside their heads with what they personally know themselves (14).
@@
Third, technological tools reduce the need for the learner to be cognitively engaged with the task at hand. For instance, knowing that we will be able to search for a fact on a computer has been shown to reduce memory for that fact, instead encouraging recall of how to search for it (15). And drivers who use GPS tend to have worse hippocampal-dependent spatial memory, both cross-sectionally and longitudinally (16, 17). To the extent that tools powered by generative AI instantaneously produce turn-key solutions for complex cognitive tasks, they may be especially detrimental to learning. It is, after all, tempting to copy and paste the output of an AI tool without even laying eyes on it.
@@
There is a plausible, albeit less obvious, alternate hypothesis. Using AI tools to help do our work could help us develop our own skills. In particular, the current generation of AI tools may teach by example, offering high-quality and personally tailored demonstrations of abstract principles that are otherwise difficult to grasp and apply. Classic research shows that worked examples of math problems (i.e., not just answers but the step-by-step process by which problems are solved) scaffold learning more effectively than explanations alone (18, 19). Compared to textbooks, conventional computer tutoring programs (e.g., (20)), and even human teachers, today’s AI tools are unprecedented in their ability to deliver hyper-personalized high-quality examples on command. Thus, AI tools may improve skill development if the benefits of exposure to just-in-time examples tailored to learners’ needs outweigh the costs of diminished engagement.
@@
There is little research on how AI tools influence skill development. In working papers, results have been mixed. Some studies have found that interacting with AI tools improves skill on subsequent tests in which AI tools are not available (21, 22), while others have shown null or even negative effects (23, 24, 22). Notably, these studies examine AI tutors, chatbots, or explanations explicitly designed to support learning, rather than simply providing solutions as is typical in real-world use. Further, they focus exclusively on mathematics and computer programming.
In this investigation, we ask whether AI tools can increase intermediate-term skill development, above and beyond just improving performance while using the AI tool. We focus on writing—the most common use of AI at work, as ranked in a nationally representative survey of American adults in August 2024 (25). Participants in our three pre-registered studies were American adults on the survey platform Prolific.
To differentiate the effects of AI use on learning versus performance, we developed a paradigm in which all participants were given a baseline writing test (i.e., revising a poorly written cover letter). They were then introduced to evidence-based strategies for professional writing, with descriptions and examples for each one (26, 27). Next, participants were randomly assigned to one of three conditions: practicing rewriting a different cover letter with access to an AI tool based on the same writing principles that users had just learned (See Figure 1), practicing without this tool, or a no-practice control group. To assess gains in writing skill, participants completed an incentivized test in which they rewrote yet another cover letter without access to AI, with a cash bonus guaranteed for submissions ranked in the top 10 percent. To assess the persistence of skill improvement, all participants were invited to complete a similar incentivized test of writing skill one day later. See Figure 2. We used GPT4o to rate each cover letter for each of the five writing principles introduced in this experiment. We averaged these ratings to produce summary scores of writing skill and, in a random subsample (n = 30), validated these scores using trained human raters (r = .83, p < .001). Additionally, we asked a separate sample of participants to read pairs of test-phase cover letters randomly selected from different conditions, and to indicate which letter would be more likely to secure a job interview. The relative likelihood of a cover letter securing an interview correlated positively with AI ratings of writing skill (rs = .29 and .28, ps < .001, for Study 2 and 3, respectively).
@@
In Study 1, forecasters presented with this design were twice as likely to predict that practicing writing with the assistance of the AI tool would impair learning compared to practicing without the AI tool. In Study 2, however, participants who had practiced with the AI tool learned more (i.e., wrote better cover letters during the test phase) compared to either comparison group—an advantage that persisted in a one-day follow-up test. Finally, in Study 3, we explored the mechanism for these learning gains by introducing an example-only condition. Participants who had merely seen an AI-generated example (but did not have an opportunity to practice) improved in writing skill as much as participants who had practiced with an AI tool; benefits again persisted in a one-day follow-up test. Test-phase cover letters written by participants who had practiced with AI (Studies 2 and 3) or had seen an AI example (Study 3) were more likely to secure hypothetical job interviews compared to cover letters written by participants who had not practiced (Study 2) or had practiced without AI (Studies 2 and 3).
Study 1
@@
We showed N = 150 participants screenshots of a random assignment study with three conditions. We asked them to rank-order these conditions according to how much they predicted future participants would learn in each. Confirming our pre-registered hypothesis, nearly twice as many forecasters (64.7%) ranked practicing alone above practicing writing with access to an AI tool as the converse (35.3%, (1) = 12.9, p < .001), see Figure 3. Participants made this prediction regardless of self-reported experience with AI (OR = 0.83, p = .239) or any other measured demographic characteristic (ps > .05).
In open-ended responses, forecasters who were pessimistic about the effect of the AI tool on learning speculated that it would crowd out effort (e.g., “Practicing alone would force more recall and problem-solving skills, while AI essentially gives the answer for them.”, “I think oftentimes using AI impedes the learning process because it’s the ‘easy way.”’). Those with positive views, on the other hand, cited the possibility of AI providing insights or examples that would be otherwise unavailable (“As much as I hate AI, I do not believe you can improve in any manner if you do not have examples or ways of learning, and AI can provide this.”)
Study 2
Study 2 tested whether the predictions of Study 1 forecasters were accurate. Specifically, N = 2,238 participants completed a baseline questionnaire and pretest (rewriting a poorly written cover letter), followed by a lesson introducing five principles of effective writing (i.e., less is more, make reading easy, design for easy navigation, use enough formatting but no more, make responding easy) (26). Next, participants were randomly assigned to one of three practice conditions: (1) rewriting a new cover letter with an AI writing tool that revised text instantly based on these principles, (2) rewriting the new cover letter without the AI tool, or (3) a no practice control. At the end of the session, all participants completed a test of writing skill (rewriting a yet another cover letter without access to the AI writing tool) and an exit questionnaire. One day later, all participants were invited to complete a similar incentivized test of writing skill.
@@
Consistent with other studies demonstrating the productivity benefits of AI tools (3, 2), participants given access to the AI writing tool produced cover letters during the practice phase that were dramatically higher in quality than participants who were not (d = 1.01, p <.001).
The learning advantage of having practiced with AI was evident in the test phase: Consistent with our pre-registered hypothesis, participants who had practiced with the AI tool produced higher-quality writing than did participants who either had practiced without the AI tool (d = .38, p <.001) or who had not practiced at all (d = .47, p <.001). See Figure S4. Likewise, cover letters written by participants who had practiced with AI were more likely to secure a hypothetical job interview than cover letters by participants who had practiced without AI (.54 vs. .50, p = .002) or had not practiced at all (.54 vs. .47, p < .001). See Figure 5. Notably, the effect of practicing with AI was not limited to stronger or weaker writers—participants across all skill levels benefited equally from practicing with AI.
@@
The learning benefits of using an AI writing tool were evident despite reduced effort during the practice phase. Compared to participants who practiced alone, participants who practiced with the AI tool spent 0.44 fewer minutes during the practice phase (3.73 vs. 4.17; d = -.12, p = .025), logged roughly a quarter as many keystrokes (26% d = -.44, p <.001) and self-reported expending less effort during practice (d = -.31, p <.001).
Nevertheless, it would be inaccurate to label writers practicing with AI as entirely disengaged. Copying, pasting, and submitting the AI tool’s output could be accomplished almost instantly. Yet, the majority of participants chose to interact with the task for over 3 minutes, and 95% made at least one edit to the AI-generated output before final submission. See Figure S5 in Supplementary Information for details.
Differential effort during practice raised the possibility that participants who had practiced with the AI tool outperformed those who had practiced on their own because they were less fatigued during the test phase. However, participants who had practiced with AI did not spend more time as those who had practiced without AI (d = .06, p = .235), but logged more keystrokes (d = .12, p = .026), and self-reported expending less effort (d = -.12, p = .023) during the test phase.
@@
Following the test phase, there were no differences by condition on self-reported knowledge or motivation to improve. Despite improving more in objectively assessed writing skill, participants who had practiced with AI reported having learned no more than those who had either practiced alone or done no practice at all (ps >.05). Self-ratings of writing skill after the practice phase were also indistinguishable between participants who practiced with AI and those who practiced without the AI tool, or in the no-practice control group (ps >.05), but participants who had practiced without AI rated their skill more highly than those who did not practice (d = .14 p = .008). Finally, compared to participants who did not practice, participants who had practiced with AI were slightly less likely to request feedback after the test phase (.65 vs. .60, p = .039), but just as likely as participants who had practiced without AI (.64 vs. .60, p = .167). See Section B4 in the Supplementary Information for details.
The effects of practicing with AI persist
To examine whether the treatment effects persisted over time, we re-contacted all participants one day later. The majority of participants responded (87%), and attrition rates did not differ by condition (13% to 14% = .68, p .710). Confirming our pre-registered hypothesis, participants who had practiced with the AI tool to practice the previous day continued to outperform those who had practiced without the tool (d = .41, p <.001) as well as those who had not practiced at all (d = .46, p <.001). Participants who had practiced without the AI tool performed no better than those who did not practice (d = .05, p = .331). See Figure 4 and Section B5 of Supplementary Information for details.
@@
Study 3
@@
To gain insight into what might be driving the benefit of practicing with AI, in Study 3 (N = 2,003), we preregistered a replication and extension in which we replaced the no-practice condition of Study 2 with an example-only condition. In this condition, participants were shown an AI-generated writing example that they could not edit. To the extent that the benefit of practicing with AI was driven by exposure to examples, the example-only condition should improve performance in the test phase as much as the practice with AI condition.
@@
As in Study 2, participants given access to the AI writing tool dramatically outperformed participants who did not get access to it, both while using it during the practice phase (d = 1.22, p <.001), and during the no-AI test phase (d = .34, p < .001). Their test-phase cover letters were also relatively more likely to secure them hypothetical job interviews (.51 vs. .47, p = .008).
Participants who had merely seen an AI-generated example also improved more in writing skill than those who had practiced without AI (d = .37, p = < .001), and produced letters that were relatively more likely to secure them interviews (.52 vs. .47, p = .007). Notably, they improved as much as participants who had practiced with the AI tool (and could edit its output, d = .03, p = .883), and were offered hypothetical interviews at similar rates as them (.51 vs. .52, p = .561). See Figures S6 and 7. As in Study 2, gains from practicing with AI and seeing AI examples were comparable across all levels of baseline writing skill.
@@
During the practice phase, participants who saw the AI example spent 2.32 fewer minutes than participants practicing with AI (d = 0.99, p < .001) and 2.96 fewer minutes than participants practicing without AI (d = 1.13 p = < .001), and reported expending less effort than those practicing with an AI tool (d = 0.19, p = .003) and those practicing without AI (d = 0.32, p < .001). As in Study 2, participants who practiced with AI logged 2.83 times fewer keystrokes than did participants who practiced without AI. As expected, participants exposed to the AI example logged 0 keystrokes.
During the test phase, participants who had seen the AI example worked for an additional 56 seconds more than participants who had practiced without AI (d = 0.22, p < .001). They also logged 30% more keystrokes than participants who had practiced with AI (d = 1.86 p <.001) and 60% more than those who practiced without AI (d = 2.39, p < .001). Across conditions, all participants self-reported similar levels of subjective effort (ds < .08, ps > .05).
@@
As in Study 2, despite learning more, participants who had practiced with AI or had merely seen an AI-generated example reported learning similar amounts to those who practiced without AI (ps > .05) and rated their writing skill after practice at comparable levels (ps > .05). Moreover, all participants requested feedback at similar rates (proportions ranged from 63% to 67%). See Section C4 in the Supplementary Information for details.
The effects of seeing an AI example persist
When we recontacted a subsample of (n = 800) participants one day later, the majority responded (n = 633, 80%); the attrition rates ranged from 17% to 24% and did not differ by condition ( (2) = 4.56, p = .102). The effect remained robust after 24 hours. Participants who had practiced with the AI-tool (d = .29, p = .006) and participants who had simply seen an AI example (d = .32, p = .003), both continued to outperform those who had practiced without the tool. Participants who had merely seen an AI example performed as well as those who had practiced with AI (d = .02, p = .830). See Figure 6 and Section C5 of Supplementary Information for details.
As in Study 2, the above findings were not moderated by individual differences. See Table S23 in the Supplementary Information for details.

### v3

# Coach not crutch: AI assistance can enhance rather than hinder skill development

@@ Abstract
Does using AI make you stupid? Contrary to predictions by lay forecasters, participants randomly assigned to practice writing with access to an AI tool improved more on a writing test one day later compared to writers assigned to practice without AI—despite exerting less effort. Likewise, whereas forecasters predicted that AI would be less helpful, and less worth paying for, than personalized feedback from experienced human editors, we found the converse: practicing with AI increased skill more than feedback from human editors (or simply googling examples and tips). The results of a third experiment suggest that AI may teach by example: writers merely shown an AI-generated cover letter (without the opportunity to practice) performed as well as writers who practiced writing with the original AI tool. Collectively, the findings of these pre-registered studies constitute an existence proof that by providing personalized, high-quality examples, AI can improve, rather than undermine, learning.
@@
Nevertheless, there is growing concern that AI tools will be used as a crutch, providing immediate gains in performance at the expense of long-run development of human capital (5, 6, 7). For instance, in a 2024 poll, 62% of surveyed adults predicted that Generative AI would “lead to humans becoming less intelligent” (8). In January 2023, New York City public schools banned ChatGPT, citing “concerns about negative impacts on student learning (9).” When this ban was lifted three months later, it was not because of AI’s potential to scaffold learning, but instead because of the “reality that students are participating in and will work in a world where understanding Generative AI is crucial (10).” However, teachers remain worried: in a nationally representative poll in May 2024, four times as many K-12 educators judged the use of AI tools as net harmful (24%) than net beneficial (6%) (11).
@@
Second, the fluent and instantaneous solutions AI tools generate may contribute to an illusion of mastery. To the extent users conflate the skills of an AI tool with their own, they may be less likely to seek feedback and improve. Prior research has found that searching for information on the Internet, for example, creates an illusion whereby people conflate knowledge outside their heads with what they personally know themselves (14).
@@
Third, technological tools reduce the learner’s need to be cognitively engaged with the task at hand. For instance, knowing that one can search for a fact on a computer has been shown to reduce memory for that fact, instead encouraging recall of how to search for it (15). And drivers who use GPS tend to have worse hippocampal-dependent spatial memory, both cross-sectionally and longitudinally (16, 17). When AI-powered tools instantaneously produce turn-key solutions for complex cognitive tasks, they may be especially detrimental to learning. It is, after all, tempting to copy and paste the output of an AI tool without even laying eyes on it.
@@
On the other hand, using AI tools to assist us in our work may, in fact, help us develop our own skills. In particular, the current generation of AI tools may teach by example, offering high-quality and personally tailored demonstrations of abstract principles that are otherwise difficult to grasp. Classic research shows that worked examples of math problems (i.e., not just answers but the step-by-step process by which problems are solved) scaffold learning more effectively than explanations alone (18, 19). It seems likely that, compared to textbooks, conventional computer tutoring programs (e.g., (20)), and human teachers, AI tools may provide higher quality, just-in-time examples exquisitely tailored to a learner’s specific needs. Indeed, it’s possible that in both quality and didactic utility, AI examples (e.g., of an excellent cover letter) might surpass even those provided by domain experts (e.g., professional writers and editors). Thus, if the upside of exposure to excellent examples tailored to the learners’ needs outweighs the downside of diminished engagement, AI tools may support skill development.
@@
The scant research on how AI tools influence skill development has yielded mixed results. Some studies report that interacting with AI tools improves skill on subsequent tests in which AI tools are not available (21, 22), while others have shown null or even negative effects (23, 24, 22). Notably, these studies examine AI tutors, chatbots, or explanations explicitly designed to support learning, rather than simply providing solutions, as is typical in real-world use. Further, they focused exclusively on mathematics and computer programming.
In this investigation, we ask whether AI tools can support skill development, above and beyond merely improving performance while using the AI tool. We focus on writing—the most common use of AI at work, as ranked in a nationally representative survey of American adults in August 2024 (26). In our experimental paradigm all participants are given a baseline writing test (i.e., revising a poorly written cover letter), followed by a short lesson on five evidence-based strategies for professional writing (25, 27). Next, participants are randomly assigned to different types of practice (e.g., practice with access to an AI tool, practice without access to an AI tool, practice with feedback from professional editors, and practice with Google Search to find examples and tips). Finally, all participants complete an incentivized test in which they rewrite cover letters without access to AI.
@@
In Study 1, forecasters presented with this design were twice as likely to predict that practicing writing with the assistance of AI would impair learning compared to practicing without it. In Study 2, however, participants who practiced with AI actually learned more (i.e., wrote better cover letters during the test phase) than participants who practiced without AI—an advantage that persisted in a one-day follow-up test. In Study 3, forecasters predicted that feedback from AI would be less helpful than feedback from professional editors and, accordingly, were less willing to pay for it. However, in Study 4, participants who practiced with AI learned more than participants who practiced with feedback from professional editors (or those who practiced with Google Search to find examples and tips). Finally, in Study 5, we explored the mechanism for these learning gains by introducing an example-only condition. Participants who saw an AI-generated example (but did not have an opportunity to practice) improved their writing as much as participants who practiced with AI; benefits again persisted in a one-day follow-up test.
Study 1: Forecasting the effect of AI on learning
We showed N = 150 participants screenshots of a random assignment study with three conditions. We asked them to rank-order these conditions according to how much they predicted future participants would learn in each.
@@
Confirming our pre-registered hypothesis, nearly twice as many forecasters (64.7%) ranked practicing alone above practicing writing with access to an AI tool as the converse (35.3%, (1) = 12.9, p < .001), see Figure 3. Participants made this prediction regardless of self-reported experience with AI (OR = 0.72, p = .096) or any other measured demographic characteristic (ps > .05). See Table S2
In open-ended responses, forecasters who were pessimistic about the effect of the AI tool on learning speculated that it would crowd out effort (e.g., “Practicing alone would force more recall and problem-solving skills, while AI essentially just gives them the answer.”, “I think oftentimes using AI impedes the learning process because it’s the ‘easy way.”’). Those with positive views, on the other hand, cited the possibility of AI providing insights or examples that would be otherwise unavailable (“As much as I hate AI, I do not believe you can improve in any manner if you do not have examples or other ways of learning, and AI can provide this.”)
Study 2: Effect of practicing with AI on learning
Study 2 tested whether the predictions of Study 1 forecasters were accurate. Specifically, N = 2,238 participants completed a baseline questionnaire and pretest (rewriting a poorly written cover letter), followed by a lesson introducing five principles of effective writing (i.e., Less is more, Make reading easy, Design for easy navigation, Use enough formatting but no more, Make responding easy) (25). Next, participants were randomly assigned to one of three practice conditions: (1) rewriting a new cover letter with an AI writing tool that instantly revised text based on these principles, (2) rewriting the new cover letter without the AI tool, or (3) a no-practice control. At the end of the session, all participants completed a test of writing skill (rewriting a new cover letter without access to the AI writing tool) and an exit questionnaire. One day later, all participants were invited to complete a similar incentivized test of writing skill. See Figure 2.
We assessed writing skill in two ways. First, we used GPT-4o to rate cover letters on each of the principles of effective writing and then averaged these ratings to produce a summary score (Cronbach’s = .81). This operationalization demonstrated convergent validity with ratings by human research assistants who rated a random subsample of n = 100 cover letters on the same principles (r = .70, p < .001). Second, we recruited participants naive to this investigation to read two letters and select the one more likely to secure its writer a job interview. These two operationalizations converged: preferred cover letters obtained higher AI ratings of writing skill (rs = .29).
@@
Consistent with other studies demonstrating the productivity benefits of AI tools (3, 2), participants given access to the AI writing tool produced cover letters during the practice phase that were dramatically higher in quality than participants without access (d = 1.01, p <.001).
The learning advantage of having practiced with AI was evident in the test phase: Consistent with our preregistered hypothesis, participants who had practiced with the AI tool produced higher-quality writing than did participants who either had practiced without the AI tool (d = .38, p <.001) or who had not practiced at all (d = .47, p <.001). See Figure S3. Likewise, cover letters written by participants who had practiced with AI were more likely to secure a hypothetical job interview than cover letters by participants who had practiced without AI (.54 vs. .50, p = .002) or had not practiced at all (.54 vs. .47, p < .001). See Figure 4B.
@@
Compared to participants who practiced alone, participants who practiced with the AI tool spent 0.44 fewer minutes during the practice phase (3.73 vs. 4.17; d = .12, p = .025, quantity of effort), logged roughly a quarter as many total keystrokes (26% d = .44, p <.001) and logged 65% fewer keystrokes per minute (d = .22, p <.001, intensity or quality of effort), and self-reported expending less effort during practice (d = .31, p <.001).
Nevertheless, it would be inaccurate to label writers practicing with AI as entirely disengaged. Rather than copying and pasting the AI tool’s output, the majority of participants practicing with AI chose to interact with the task for over 3 minutes, and 95% made at least one edit to the AI-generated output before final submission. See Figure S4 in Supplementary Information for details.
@@
In the exit questionnaire, participants who had practiced with AI reported learning about as much as participants who had practiced without it or not practiced. Likewise, they judged their writing skill similarly to those who had either practiced alone or done no practice at all (|ds| 0.10, ps >.05). When asked if they wanted to see additional feedback, they were about as likely to say yes as participants who had practiced without AI (.64 vs. .60, p = .167), and slightly less likely than those who had not practice at all (.65 vs. .60, p = .039). See Section C4 in the Supplementary Information for details.
The benefits of practicing with AI were just as large a day later
To examine whether the treatment effects persisted over time, we re-contacted all participants one day later. Most participants responded (87%), and attrition rates did not differ by condition (13% to 14% = .68, p = .710). Confirming our pre-registered hypothesis, participants who had practiced with the AI tool the previous day continued to outperform those who had practiced without the tool (d = .41, p <.001) as well as those who had not practiced at all (d = .46, p < .001). Participants who had practiced without the AI tool performed no better than those who did not practice (d = .05, p = .331). The magnitude of the follow-up effect size was not significantly different from the immediate test phase effect size (condition time interactions were non-significant (ps > .69). See Figure 4A and Section C5 of Supplementary Information for details.
Results were not moderated by individual differences
@@
Study 3: Willingness to pay for AI vs. human editors
Given unwarranted pessimism about the pedagogical power of AI, in Study 3 we examined whether people underinvest in AI as a learning resource. Specifically, we asked N = 150 people to imagine they were participants in Study 2 and asked them how much they would be willing to pay for feedback from the AI tool vs. feedback from experienced human editors.
Participants were willing to pay more for feedback from expert human editors than for AI feedback
Out of a maximum of a dollar (the total amount of the hypothetical bonus participants could earn), participants on average were willing to pay 28 cents for the AI tool vs. 38 cents for feedback from an experienced human editor (t(149) = 4.83, p <.001, d = 0.39).
Participants predicted that editorial feedback would be more beneficial
Participants predicted that experienced human editors (M = 6.68 on a 0—10 Likert scale) would be more effective in helping them improve their writing, compared to feedback from an AI tool (M = 5.64; t(148) = 4.77, p <.001, d = 0.39).
Predicted effectiveness mediated underinvestment in AI feedback
People’s beliefs about how effective AI and human editor feedback would be predicted willingness to pay (b = 0.06, SE = 0.01, p < .001) and accounted for most of the preference for human editors (indirect effect = 0.06, p < .001). Once these predictions were accounted for, the remaining preference for human feedback was reduced by 60% (c’ = 0.04, SE = 0.02, p = .054). See Figure 5 and Section D1 of Supplementary Information for details.
Study 4: Practicing with AI vs. with expert feedback (vs. with Google Search)
Participants (N = 2,997) completed a baseline questionnaire and pretest, followed by the same lesson on effective writing as in Study 2, and then practiced rewriting a new cover letter on their own. Next, participants were randomly assigned to one of three conditions: (1) practice with AI, (2) practice with feedback from professional human editors, (3) practice with Google Search. Letters written by participants assigned to the second condition were given to professional editors averaging 25 years of professional editorial and writing experience. These 49 editors, recruited in partnership with the Journalists’ Resource at Harvard Kennedy School, provided personalized revisions and constructive feedback. The following day, all participants attempted to improve the cover letter they had previously submitted, using the resources afforded by their assigned condition. Finally, all participants completed a writing test (rewriting a cover letter without access to AI) and an exit questionnaire.
Practicing with AI improved writing skill more than feedback from professional editors and Google Search
During the practice phase, participants working with AI outperformed participants who received personalized feedback from professional editors (d = 0.76, p <.001) and those assigned to Google cover letter examples and tips (d = 1.03, p <.001).
During the test phase, in which participants were asked to write a new cover letter without access to any outside resources, participants who had practiced with the AI tool produced higher-quality cover letters than those who had practiced with Google Search (d = 0.46, p <.001), and even outperformed those who had practiced with personalized feedback from professional editors (d = 0.20, p <.001). See Figure 6A.
Cover letters written by participants who had practiced with AI were as likely to secure a hypothetical job interview as cover letters by participants who had practiced with editor feedback (.51 vs. .51, p = .843), and more likely to than those written by participants who had practiced with Google Search (.51 vs. .47, p = .024). See Figure 6B.
A possible concern is that reviewing many letters consecutively caused editors to get tired and give worse feedback over time. This could suggest that our experiment compared AI against a “strawman” version of human expertise. However, analysis of the editors’ performance revealed the opposite pattern: while editors became more efficient, spending less time per letter as the task progressed (b = 0.387, p < .001), the quality of their edits and feedback actually improved (b = 0.013, p < 0.001). Thus, the AI tool’s advantage over professional editors cannot be attributed to a decline in editorial performance. See Section E2 of Supplementary Information for details.
Practicing with AI was not more effortful than getting feedback from professional editors and was less effortful than looking for examples online
During the practice phase, participants with access to AI spent about the same time as those who practiced with editor feedback (difference = 0.43 minutes, d = 0.09, p > .05), but about 1.18 fewer minutes than those who practiced with Google Search (d = 0.32, p < .001). Nevertheless, participants across all three groups reported similar levels of subjective effort (ps > .05). Likewise, participants who practiced with AI logged about as many keystroles as those who practiced with editor feedback (68% fewer, d = 0.08, p > .05) and 76% fewer keystrokes than those who practiced with Google Search (d = 0.32, p < .001). They also typed at a slower rate than participants who practiced with editor feedback (d = 0.41, p < .001) or those who practiced with Google Search (d = 0.34, p < .001).
No evidence of AI-induced illusion of mastery
Participants who had practiced with AI did not report learning more than those who had practiced with editor feedback (d = 0.07, p > .05), nor did they consider their writing skills to be better (d = 0.02, p > .05). Both groups, however, correctly recognized that they had learned more than those who had practiced with Google Search (ds = .28 and .20, respectively), and evaluated their writing skills accordingly (ds = .11 and .12, respectively). Participants who had practiced with AI and those who had practiced with editor feedback requested feedback on their test submissions at similar rates (71% and 70%, respectively). Both groups were more likely to request feedback than those who had practiced with Google search (64%, p = .003 and .007, respectively). See Section E6 in the Supplementary Information for details.
Results were not moderated by individual differences
As in Study 2, the above findings were consistent across all measured individual differences. See Table S22 in the Supplementary Information for details.
Study 5: Examples as the mechanism for AI learning gains
@@
To better understand what drives the benefit of AI practice, in Study 5 (N = 2,003), we preregistered a replication and extension in which we replaced the no-practice condition of Study 2 with an example-only condition. In this condition, we showed participants an AI-generated writing example that they could not edit. To the extent that the benefit of practicing with AI was driven by exposure to examples, the example-only condition should improve performance in the test phase as much as the practice with AI condition.
@@
As in Study 2, participants given access to the AI writing tool dramatically outperformed participants who did not get access to it, both while using it during the practice phase (d = 1.22, p <.001), and during the no-AI test phase (d = .34, p < .001). Their test phase cover letters were also relatively more likely to secure them hypothetical job interviews (.51 vs. .47, p = .008).
Participants who had merely seen an AI-generated example also improved more in writing skill than those who had practiced without AI (d = .37, p < .001), and produced letters that were relatively more likely to secure them interviews (.52 vs. .47, p = .007). Notably, they improved as much as participants who had practiced with the AI tool (and could edit its output, d = .03, p = .883), and were offered hypothetical interviews at similar rates as them (.51 vs. .52, p = .561). See Figures S10 and 7B.
@@
During the practice phase, participants who saw the AI example spent 2.32 fewer minutes than participants practicing with AI (d = 0.99, p < .001) and 2.96 fewer minutes than participants practicing without AI (d = 1.13 p = < .001), and reported expending less effort than those practicing with an AI tool (d = 0.19, p = .003) and those practicing without AI (d = 0.32, p < .001). As in Study 2, participants who practiced with AI logged 74% fewer keystrokes and 68% fewer keystrokes per minute (d = 0.15, p = .007) compared to participants who practiced without AI. As expected, participants exposed to the AI example logged 0 keystrokes.
@@
As in Studies 2 and 4, despite learning more, participants who had practiced with AI or had merely seen an AI-generated example reported learning similar amounts to those who practiced without AI (ps > .05) and rated their writing skill after practice at comparable levels (ps > .05). Moreover, all participants requested feedback at similar rates (proportions ranged from 63% to 67%). See Section F4 in the Supplementary Information for details.
The benefits of seeing an AI example were just as large a day later
When we recontacted a subsample of participants (n = 800) one day later, the majority responded (n = 633, 80%); the attrition rates ranged from 17% to 24% and did not differ by condition ( (2) = 4.56, p = .102). The effect remained robust after 24 hours. Participants who had practiced with the AI-tool (d = .29, p = .006) and participants who had simply seen an AI example (d = .32, p = .003), both continued to outperform those who had practiced without the tool. Participants who had merely seen an AI example performed as well as those who had practiced with AI (d = .02, p = .830). As in Study 2, the magnitude of the follow-up effect size was not significantly different from the immediate test phase effect size (condition time interactions were non-significant (ps > .96). See Figure 7A and Section F5 of Supplementary Information for details.
Results were not moderated by individual differences
As in the previous studies, the above findings were consistent across all measured individual differences. See Table S35 in the Supplementary Information for details.